OpenStudy (anonymous):

Given: f(x) = x - 7 and h(x) = 2x + 3 Write the rule for h(f(x)). h(f(x)) = 2x - 11

OpenStudy (anonymous):

h(f(x)) = 2x - 11 h(f(x)) = 2x - 7 h(f(x)) = 2x - 4 h(f(x)) = 3x - 4

jhonyy9 (jhonyy9):

just substitute the f(x) in place of x inside equation h(x) so f(x)=x-7 h(x)= 2x+3 so h(f(x)) = 2(x-7) +3 = 2x-14+3 = 2x-11 h(f(x)) = 2x-11 hope this is understandably sure easy